If n=16, sample mean = 75, sample standard deviation s = 8. The hypotheses are Ho: μ = 80 and Ha: μ <80. What is the test statistic? Please show your work.
Answer by Boreal(15235)

Ho is mu > =80
Ha is mu <80. Must have the entire possibility contained by both.
One way test.
test statistic is t df=15= (75-80)/8/sqrt (16); (data-hypothesized mean)/s/sqrt (n). It is a t df=(n-1)
The critical value (not asked for) is =-20/8=-2.5